# RED-PSM: Regularization by Denoising of Partially Separable Models for Dynamic Imaging

### v3

@@ VI-D3 Effect of Initialization
Combined also with the convergent algorithm eliminating the need for a unrealistic stopping oracle and the theoretical analysis, RED-PSM provides improved reliability which is of practical significance.
Figure 9: PSNR vs. iteration comparison for RED-PSM as described in Section VI-C. The area between the best and the worst PSNR for each iteration is shown in blue to highlight the varying performances of five different runs with random initialization.
@@ VI-D6 Patch-based RED denoiser
[…]
To showcase the potential of the suggested scheme, we replace the full-size $D_{\phi}$ with a patch-based counterpart in the RED step and compare the performance with the originally proposed method for 2D dynamic objects.
@@
Table IV compares the results for the two denoiser types for both objects, using the same denoiser training policy as in Section VI-C, and experimental configurations of Table VI in the Supplementary Material Section A-A. The results show little difference, thus verifying the effectiveness of the patch-based version. The analysis of the computational requirements of the patch-based RED-PSM variant in Section IV-D shows its potential for a highly scalable implementation.
We also note that in a divergent beam scenario, the contribution of a patch to a projection will be position dependent and this would require accurate bookkeeping. To this end, there are methods on tile-based approaches in fan-beam [89] and cone-beam [90] tomography. However, we leave such analysis for future studies.
@@ VII Conclusions
We proposed RED-PSM, the first PSM-based approach to dynamic imaging using a pre-trained and learned (RED-based) spatial prior. The objective in the proposed variational formulation is optimized using a novel and effective bilinear ADMM algorithm, which enforces the PSM as a hard constraint. Unlike existing PSM-based techniques, RED-PSM is supported by theoretical analysis, with a convergence guarantee to a stationary point of the objective. The results of the numerical experiments show better reconstruction accuracy and considerably faster run times compared to a recent DIP-based algorithm. A patch-based regularizer version of RED-PSM provides almost equivalent performance with massive reduction of storage requirements, indicating the potential of our framework for dynamic high-resolution 2D or 3D settings.
@@
Appendix A Proof of Theorem 1
@@ 2. There exists an $\underline{\mathcal{L}_{\beta}}$ such that
where the first equality in (23) follows from the dual variable update in Step (5) of Algorithm 1. The last two terms on the RHS of (A) are further split into
@@
$\quad+\mathcal{L}_{\beta}[\Psi^{(i-1)},\Lambda^{(i)},f^{(i-1)};\gamma^{(i-1)}]-\mathcal{L}_{\beta}[W^{(i-1)}].$ (24)
The first two terms on the RHS of (A) are upper bounded using the strong convexity of $S_{f}$ with modulus $\alpha_{f}$:
@@
where the last line of (A) is due to the optimality condition of Line 4 of Algorithm 1 for $f^{(i)}$.
The next two terms on the RHS of (A) are upper bounded using the strong convexity of $S_{\Psi}$ with modulus $\alpha_{\Psi}$ as
@@
Repeating similar steps for the last two terms on the RHS of (A) using the objective $S_{\Lambda}$ leads to the upper bound of $-\frac{\alpha_{\Lambda}}{2}\|\Lambda^{(i)}-\Lambda^{(i-1)}\|_{F}^{2}$.
@@
where (A) follows by using (20) to replace the second term of (26) by that of (A) and assuming $\beta>L$, and (28) is due to the gradient Lipschitz continuity of the regularizer $\bar{\rho}(\cdot)$,
@@
It follows from Lemma 6 that the gap $L_{\beta}[W^{(i)}]-H(\Lambda^{(i)}\Psi^{(i)T},\Lambda^{(i)},\Psi^{(i)})$ between the augmented Lagrangian and the objective converges to zero, and therefore thanks to the convergence of $L_{\beta}[W^{(i)}]$, the objective $H$ too converges. This establishes Part (i) of Theorem 1.
@@
It follows that the sequences of iterates $\{(f^{(i)},\Lambda^{(i)},\Psi^{(i)},\gamma^{(i)})\}$ has at least one accumulation point $(f^{*},\Lambda^{*},\Psi^{*},\gamma^{*})$. Next, we establish the properties of all such accumulation points.
@@
Supplementary Material
A-A Experimental configurations
(a) Walnut (b) Comp. Material
32 PSM-TV-S (R) 3 4 5e-2 - - 3 4 10 - -
32 PSM-TV-ST (R) 4 5 5e-2 5e-2 - 3 4 1e1 1e2 -
32 PSM-RED (P) 3 7 1e-4 - 1e-4 3 9 5.12e-2 - 1.6e-2
64 PSM-TV-S (R) 3 4 5e-2 - - 5 6 10 - -
64 PSM-TV-ST (R) 4 5 5e-2 5e-2 - 5 6 1e1 1e2 -
64 PSM-RED (P) 4 7 1e-4 - 5e-4 5 9 32e-4 - 4e-3
128 PSM-TV-S (R) 5 7 5e-2 - - 8 9 5 - -
128 PSM-TV-ST (R) 6 7 5e-2 1e-1 - 8 9 5 5e1 -
128 PSM-RED (P) 6 9 2e-4 - 2e-4 8 11 4e-4 - 2e-3
256 PSM-TV-S (R) 10 11 5e-2 - - 11 13 10 - -
256 PSM-TV-ST (R) 10 11 5e-2 5e-2 - 11 13 1e1 1e2 -
256 PSM-RED (P) 10 11 5e-5 - 1e-4 11 13 1e-4 - 1e-3
(a) Walnut (b) Comp. Material
$P$ Denoiser $K$ $d$ $\lambda$ $\beta$ $K$ $d$ $\lambda$ $\beta$
256 Full-image 10 11 5e-5 1e-4 11 13 1e-4 1e-3
256 Patch-based 10 11 5e-5 1e-4 11 13 1e-4 1e-3
Dataset ##\## of layers ##\## of channels Denoising
Walnut 6 64 Direct
Compressed material 3 32 Residual
Cardiac dMRI 6 64 Residual
TABLE VII: Denoiser DnCNN configurations for different datasets.
A-B Time and space complexity analyses
A-B1 Storage requirements for the input quantities
A-B2 RED-PSM
Expanding the $\Lambda$ and $\Psi$ subproblems in Algorithm 2 in $t$, we have
$+\xi\|\Lambda\|_{F}^{2},$ (32)
and
$+{\beta}\|(\Lambda\Psi^{T}-f+\gamma)e_{t}\|_{2}^{2}.$ (33)
Starting with the gradient of (A-B2) with respect to $\Psi^{T}e_{t}$, we have
$+2\Lambda^{T}(\gamma-f)e_{t}.$ (34)
When $\Psi=UZ$ is used, the gradient for $Z$ requires the multiplication of (A-B2) with $U^{T}e_{t}$, adding a minor operation count $O(dK)$.
Next, we consider the gradient of (A-B2) with respect to $\Lambda$
$+(R_{\theta(t)}^{T}g_{t}+(\gamma-f)e_{t})(e_{t}^{T}\Psi).$ (35)
A-B3 Patch-based RED-PSM
A-C Sample reconstructed spatial and temporal basis functions
To help interpret the operation of the proposed RED-PSM algorithm, we display the reconstructed spatial and temporal basis functions $\Lambda$ and $\Psi$ for the time-varying walnut setting with $P=256$ and DCT-II latent temporal basis $U$ in Figure 12. The energy corresponding to each basis pair $E_{k}=\|\Lambda_{k}\Psi_{k}^{T}\|_{F}^{2}$ is shown in Figure 13.
Figure 13: $E_{k}=\|\Lambda_{k}\Psi_{k}^{T}\|_{F}^{2}$ for each spatiotemporal basis pair for the setting described in Section A-C.

### v4

@@ VI-D3 Effect of Initialization
Combined also with the convergent algorithm eliminating the need for an unrealistic stopping oracle and the theoretical analysis, RED-PSM provides improved reliability which is of practical significance.
Figure 9: Advantage of ProSep-based vs. random initialization of RED-PSM (see Sec. VI-C). For random initialization, the area in blue between the best and the worst PSNR for each iteration highlights the varying performances of five different runs with different random initializations.
@@ VI-D6 Patch-based RED denoiser
To improve the scalability of the method to higher resolution and/or 3D dynamic object, conveniently, the objective in (5) can be manipulated to operate on the patches of temporal image frames of the time-varying object. This circumvents the need to store the complete image frame at a given time, and also enables the denoiser $D_{\phi}$ to be both trained and to operate on patches of image frames. To showcase the potential of the suggested scheme, we replace the full-size $D_{\phi}$ with a patch-based counterpart in the RED step and compare the performance with the originally proposed method for 2D dynamic objects.
@@
Table IV compares the results for the two denoiser types for both objects, using the same denoiser training policy as in Section VI-C, and experimental configurations of Table VI in the Supplementary Material Section A. The results show little difference between the patch-based denoiser and full frame-based denoiser RED-PSM variants, thus verifying the effectiveness of the patch-based version. The analysis of the computational requirements of the patch-based RED-PSM variant in Section IV-D shows its potential for a highly scalable implementation.
We note that in a divergent beam scenario, the contribution of a patch to a projection will be position-dependent and this would require accurate bookkeeping. Details for doing so can be found in tile-based methods for fan-beam [139] and cone-beam [140] tomography. However, we leave such analysis for future studies.
@@ VII Conclusions
We proposed RED-PSM, the first PSM-based approach to dynamic imaging using a pre-trained and learned (RED-based) spatial prior. The objective in the proposed variational formulation is optimized using a novel and effective bi-convex ADMM algorithm, which enforces the PSM as a hard constraint. Unlike existing PSM-based techniques, RED-PSM is supported by theoretical analysis, with a convergence guarantee to a stationary point of the objective. The results of the numerical experiments show better reconstruction accuracy and considerably faster run times compared to a recent DIP-based algorithm. A patch-based regularizer version of RED-PSM provides almost equivalent performance with a massive reduction of storage requirements, indicating the potential of our framework for dynamic high-resolution 2D or 3D settings.
@@
Supplementary Material
VII-A Experimental configurations
(a) Walnut (b) Comp. Material
32 PSM-TV-S (R) 3 4 5e-2 - - 3 4 10 - -
32 PSM-TV-ST (R) 4 5 5e-2 5e-2 - 3 4 1e1 1e2 -
32 PSM-RED (P) 3 7 1e-4 - 1e-4 3 9 5.12e-2 - 1.6e-2
64 PSM-TV-S (R) 3 4 5e-2 - - 5 6 10 - -
64 PSM-TV-ST (R) 4 5 5e-2 5e-2 - 5 6 1e1 1e2 -
64 PSM-RED (P) 4 7 1e-4 - 5e-4 5 9 32e-4 - 4e-3
128 PSM-TV-S (R) 5 7 5e-2 - - 8 9 5 - -
128 PSM-TV-ST (R) 6 7 5e-2 1e-1 - 8 9 5 5e1 -
128 PSM-RED (P) 6 9 2e-4 - 2e-4 8 11 4e-4 - 2e-3
256 PSM-TV-S (R) 10 11 5e-2 - - 11 13 10 - -
256 PSM-TV-ST (R) 10 11 5e-2 5e-2 - 11 13 1e1 1e2 -
256 PSM-RED (P) 10 11 5e-5 - 1e-4 11 13 1e-4 - 1e-3
(a) Walnut (b) Comp. Material
$P$ Denoiser $K$ $d$ $\lambda$ $\beta$ $K$ $d$ $\lambda$ $\beta$
256 Full-image 10 11 5e-5 1e-4 11 13 1e-4 1e-3
256 Patch-based 10 11 5e-5 1e-4 11 13 1e-4 1e-3
Dataset ##\## of layers ##\## of channels Denoising
Walnut 6 64 Direct
Compressed material 3 32 Residual
Cardiac dMRI 6 64 Residual
TABLE VII: Denoiser DnCNN configurations for different datasets.
VII-B Time and space complexity analyses
VII-B1 Storage requirements for the input quantities
VII-B2 RED-PSM
Expanding the $\Lambda$ and $\Psi$ subproblems in Algorithm 2 in $t$, we have
$+\xi\|\Lambda\|_{F}^{2},$ (18)
and
$+{\beta}\|(\Lambda\Psi^{T}-f+\gamma)e_{t}\|_{2}^{2}.$ (19)
Starting with the gradient of (VII-B2) with respect to $\Psi^{T}e_{t}$, we have
$+2\Lambda^{T}(\gamma-f)e_{t}.$ (20)
When $\Psi=UZ$ is used, the gradient for $Z$ requires the multiplication of (VII-B2) with $U^{T}e_{t}$, adding a minor operation count $O(dK)$.
Next, we consider the gradient of (VII-B2) with respect to $\Lambda$
$+(R_{\theta(t)}^{T}g_{t}+(\gamma-f)e_{t})(e_{t}^{T}\Psi).$ (21)
VII-B3 Patch-based RED-PSM
VII-C Sample reconstructed spatial and temporal basis functions
To help interpret the operation of the proposed RED-PSM algorithm, we display the reconstructed spatial and temporal basis functions $\Lambda$ and $\Psi$ for the time-varying walnut scenario with $P=256$ and DCT-II latent temporal basis $U$ in Figure 12. The energy corresponding to each basis pair $E_{k}=\|\Lambda_{k}\Psi_{k}^{T}\|_{F}^{2}$ is shown in Figure 13.
Figure 13: $E_{k}=\|\Lambda_{k}\Psi_{k}^{T}\|_{F}^{2}$ for each spatiotemporal basis pair for the setting described in Section VII-C.
VII-D Proof of Theorem 1
@@ 2. There exists an $\underline{\mathcal{L}_{\beta}}$ such that
where the first equality in (27) follows from the dual variable update in Step (5) of Algorithm 1. The last two terms on the RHS of (VII-D) are further split into
@@
$\quad+\mathcal{L}_{\beta}[\Psi^{(i-1)},\Lambda^{(i)},f^{(i-1)};\gamma^{(i-1)}]-\mathcal{L}_{\beta}[W^{(i-1)}]$ (28)
The first two terms on the RHS of (VII-D) are upper bounded using the strong convexity of $S_{f}$ with modulus $\alpha_{f}$:
@@
where the last line of (VII-D) is due to the optimality condition of Line 4 of Algorithm 1 for $f^{(i)}$.
The next two terms on the RHS of (VII-D) are upper bounded using the strong convexity of $S_{\Psi}$ with modulus $\alpha_{\Psi}$ as
@@
Repeating similar steps for the last two terms on the RHS of (VII-D) using the objective $S_{\Lambda}$ leads to the upper bound of $-\frac{\alpha_{\Lambda}}{2}\|\Lambda^{(i)}-\Lambda^{(i-1)}\|_{F}^{2}$.
@@
where (VII-D) follows by using (24) to replace the second term of (30) by that of (VII-D) and assuming $\beta>L$, and (32) is due to the gradient Lipschitz continuity of the regularizer $\bar{\rho}(\cdot)$,
@@
It follows from Lemma 6 that the gap $L_{\beta}[W^{(i)}]-H(\Lambda^{(i)}\Psi^{(i)T},\Lambda^{(i)},\Psi^{(i)})$ between the augmented Lagrangian and the objective converges to zero, and therefore thanks to the convergence of $L_{\beta}[W^{(i)}]$, the objective $H$ converges too. This establishes Part (i) of Theorem 1.
@@
It follows that the sequences of iterates $\{f^{(i)},\Lambda^{(i)},\Psi^{(i)},\gamma^{(i)})\}$ has at least one accumulation point $(f^{*},\Lambda^{*},\Psi^{*},\gamma^{*})$. Next, we establish the properties of all such accumulation points.